find the general solution of the following equation
sinx+cosx=1

Answer by lwsshak3(11628)

find the general solution of the following equation
For:(0,90˚)
sinx+cosx=1
sinx+√(1-sin^2x)=1
√(1-sin^2x)=1-sinx
square both sides
1-sin^2x=1-2sinx+sin^2x
2sin^2x-2sinx=0
2sinx(2sinx-1)=0
2sinx=0
x=0
or
sinx=1/2
x=30˚(reject, extraneous root)